If you buy a home with less than 20% down, you will pay an additional monthly fee, PMI (private mortgage insurance), until you reach 80% equity. Keep track of when you reach 80% equity so you can request to have your PMI removed. Ken Buckmiller’s home recently appraised at $292,000. His mortgage was for $277,000 at 5% for 30 years with PMI of $249.17 per month. What is his monthly payment plus PMI? His mortgage balance is currently $170,500. Has he reached 80% equity? (Use Table 15.1.) (Do not round intermediate calculations. Round your answer to the nearest cent.)
Answer #1
Monthly payment = (P*R*(1+R)^N)/((1+R)^N-1)
a P*R (277000)*(.05/12) = 1154.17
b (1+R)^N ((1+(0.05/12))^360) 4.4677
c (1+R)^N-1 ((1+(0.05/12))^360)-1 3.4677
d b/c 4.4677/3.4677 1.2884
e a*e 1154.17*1.2884 1487.03
N = 12*30 = 360
P = 277000
R = Monthly interest rate = .05/12
Additional monthly fee = $249.17
Monthly payment = $1487.03
Total monthly payment = $249.17+ $1487.03 = $1736.2
Mortgage value / market value of home = 170500/292000
58.39%
Current mortgage value balance is 170500 which is 58.39% of market value of home.
It describes that now the equity value is 41.61 which is higher than 20% equity
requirement so it could be better to removed additional fee of $249.17